40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 325     Entry time: Wed Feb 20 11:34:17 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: laser head temp is up 
MOPA head temp is running at 20.3C now
Nomally it is at 18.5C
Attachment 1: htempup.jpg  89 kB  | Hide | Hide all
htempup.jpg
ELOG V3.1.3-